OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

do you use the Sn=n/2(a1+an) that's what I am doing but i cant get the right answer....

OpenStudy (anonymous):

is it 198 and 1/3?

OpenStudy (mendicant_bias):

@cerickson2 , you are using a sum formula to attempt to find a specific value of the sequence. What you want to do in order to find this out is to use a different formula, one moment.

OpenStudy (anonymous):

How would you do that?

OpenStudy (anonymous):

@msasu25 the answer choices are 203 197 168 163

OpenStudy (mendicant_bias):

The formula for finding any term of a given arithmetic sequence is: \[a_{n}=a_{1}+(n-1)d\]What you need to do, first off, is use your two given term and given n to solve for d. Use the term wth n_18 as your a-sub-n, and use the term with the lower n as your a-sub-1.

OpenStudy (anonymous):

what numbers are used in which places?

OpenStudy (mendicant_bias):

\[a_{n}=a_{1}+(n-1)d, \ \text{where}\]\[a_{n}=a_{18}=95; \ \ \ a_{1}=-7; \ \ \ n=18; \ \ \ d= \ ?\]

OpenStudy (anonymous):

so it would look like a18 = -7 +(18 - 1) d ??????

OpenStudy (mendicant_bias):

Correct! Solve for n, when you have 95 on the other side of the equation.

OpenStudy (anonymous):

wait, where would 95 go?

OpenStudy (anonymous):

a18=-7+(18-1)d a18=-7+(17)d a18=10d I dont know what to do next

OpenStudy (anonymous):

I don't think you can add the -7 to the 17d

OpenStudy (anonymous):

oh... hah then what would i do?

OpenStudy (anonymous):

I have no idea lol

OpenStudy (anonymous):

thats ok @Mendicant_Bias

OpenStudy (mendicant_bias):

I think you don't understand how equations work, one moment.

OpenStudy (anonymous):

I mean the 95 would go on the lhs then add 7 to both sides

OpenStudy (anonymous):

ok I got the answer now

OpenStudy (mendicant_bias):

@cerickson2 , could you take a shot at writing out what the equation looks like to you, without any values? Just variables?

OpenStudy (anonymous):

what do you mean...?

OpenStudy (anonymous):

like in words, say nth term = first term + (blahl balh)

OpenStudy (mendicant_bias):

Not even in words, with symbols.

OpenStudy (anonymous):

actually first write out an arithmetic sequence 3 + 5 + 7 + 9 + 11

OpenStudy (mendicant_bias):

I gave you an equation, can you write it down, without numbers, again?

OpenStudy (anonymous):

no idea how....

OpenStudy (mendicant_bias):

You use symbols, like this: a_1+a_2=n+d+blahblahblah

OpenStudy (anonymous):

do you know the definition of an arithmetic sequence?

OpenStudy (anonymous):

no

OpenStudy (mendicant_bias):

I used symbols, to describe a relationship between numbers. That's what an equation is, cerickson.

OpenStudy (anonymous):

ok you need to start there

OpenStudy (mendicant_bias):

Hold off on the arithmetic sequence stuff. He needs to at least understand what functions and equations are first.

OpenStudy (anonymous):

provide an example of an arithmetic sequence

OpenStudy (anonymous):

i guess, ok gl guys

OpenStudy (mendicant_bias):

But yeah, make me a fictional equation, @cerickson2 , just make one up, doesn't matter what it is. Can you do that? Just make a relationship between a bunch of symbols.

OpenStudy (anonymous):

can you give me an example cause you lost me a long time ago...

OpenStudy (mendicant_bias):

\[5+a=b\]

OpenStudy (mendicant_bias):

That is an example of an equation. That is what an equation is.

OpenStudy (anonymous):

can you just make up a problem just like the one I have a question on. All I need is to see how to do a similar problem and then Ill figure it out...

OpenStudy (mendicant_bias):

I'll do that, but there's a few things you have to do first. I give you my word you'll both get your answer and understand how it works. Could you make up an equation and give me an example of it? An equation is a relation between a bunch of things that has an equals sign in the middle. Can you do that?

OpenStudy (anonymous):

like 35+1*x=72 ....like that?

OpenStudy (mendicant_bias):

Exactly! Excellent, awesome! That is an equation. Great. Alright, so I gave you an equation that describes how an arithmetic series works, but @msasu25 had a good idea in that he asked you to try to write out an arithmetic sequence, could you do that? It's okay if you can't, but take a shot at it, and if not, I'll show you how.

OpenStudy (anonymous):

yeah... I wouldnt know how to do that

OpenStudy (mendicant_bias):

Alright, no problem. Here's some examples of arithmetic series: \[S = 2, \ 4,\ 6,\ 8,\ 10,\ \dots\] \[S = 1, \ 3, \ 5, \ 7, \ 9 \dots\] \[S=5, \ 10, \ 15, \ 20, \ 25, \ \dots\]

OpenStudy (mendicant_bias):

Alright, I'll be right back, but I have to go.

OpenStudy (mendicant_bias):

~15 minutes.

OpenStudy (anonymous):

whats the answer to my question though.... i dont have time... Ill find someone else.......

OpenStudy (mendicant_bias):

I have to go to an advising appointment, if you're asking for answers for a test or homework that is due today, I cannot help you other than teaching you how to do it and helping you get to the answer. Actually, I'll give it to you, as long as you come back and learn how to do it, okay?

OpenStudy (mendicant_bias):

Does that sound good, @cerickson2 ?

OpenStudy (anonymous):

yes! I would love to learn how to do it

OpenStudy (mendicant_bias):

Alright, your answer is 197.

OpenStudy (mendicant_bias):

Good luck, and I'll be back later.

OpenStudy (anonymous):

thank you. Ill keep the question open so when you get back you can teach me

OpenStudy (anonymous):

file://localhost/Users/sylvanerickson/Pictures/Photo%20Booth/Photo%20on%202014-12-11%20at%2010.42.jpg

OpenStudy (anonymous):

its a different problem but my boyfriend helped me.

OpenStudy (mendicant_bias):

Your file is hosted on your computer, I can't connect to that, but sure thing, it's good that you figured it out!